# Beyond Grids: Multi-objective Bayesian Optimization With Adaptive Discretization

### v3

@@ Abstract
[…]
@@ 1 Introduction
We prove information-type and metric dimension-type upper bounds on the sample complexity of Adaptive $\bm{\epsilon}$-PAL (Theorem 1). Our information-type bound yields a sample complexity upper bound of $\tilde{O}(g(\epsilon))$ where $\epsilon=\min_{j}\epsilon^{j}$, $g(\epsilon)=\min\{\tau\geq 1:\sqrt{\gamma_{\tau}/\tau}\leq\epsilon\}$, and $\gamma_{\tau}$ is the maximum information gain after $\tau$ evaluations. To the best of our knowledge, this is the first information type bound for dependent objectives in the context of PAL. In addition, our metric dimension-type bound yields a sample complexity of $\tilde{O}(\epsilon^{-(\frac{\bar{D}}{\alpha}+2)})$ for all $\bar{D}>D_{1}$, where $D_{1}$ represents the metric dimension of $({\cal X},d)$ and $\alpha\in(0,1]$ represents the Hölder exponent of the metric induced by the GP on ${\cal X}$. As far as we know, this is the first metric dimension-type bound in the context of PAL. Our bounds complement each other: as we show in the appendix, neither of them dominates the other for all possible GP kernels. Specifically, we provide an example (Proposition 1) under which the information-type bound can be very loose compared to the metric dimension-type bound. Besides theory, we also show via simulations on multi-objective functions that Adaptive $\bm{\epsilon}$-PAL significantly improves over its competitors in terms of accuracy.
Organization. We provide a detailed comparison with related work in Section 2. We explain the properties of the function to be optimized and the structure of the design space in Section 3. This is followed by the description of Adaptive $\bm{\epsilon}$-PAL in Section 4. We give sample complexity bounds for Adaptive $\bm{\epsilon}$-PAL in Section 5, discuss the main aspects of computational complexity analysis in Section 6. We devote Section 7 to the computational experiments, followed by conclusions in Section 8. We present the proof of the main theorem separately in Section A, the appendix. At the end of the paper, we include a table for the frequently used notation.
@@ 2.1 Multi-objective optimization
Auer et al. (2016) consider a finite set of designs and formulate the identification of the Pareto front as a pure exploration multi-armed bandit (MAB) problem in the fixed confidence setting. Under the assumption that the centered outcomes are $1$-subGaussian and independent, they provide gap-dependent bounds on its sample complexity, which, in the single objective case, yield the well-known gap-dependent sample-complexity bounds for pure exploration in the multi-armed bandit setting (Mannor & Tsitsiklis, 2004). Other works on Multi-objective MAB include (Xu & Klabjan, 2023; Busa-Fekete et al., 2017; Öner et al., 2018). As opposed to their frequentist approach, our approach is Bayesian, which imposes a Gaussian process prior to the latent function of interest.
@@
Ours Compact A GP sample Yes Inf. & dim.-type Used
@@ 3 Background and formulation
Throughout the paper, let us fix a positive integer $m\geq 2$ and a compact metric space $(\mathcal{X},d)$. We denote by $\mathbb{R}^{m}$ the $m$-dimensional Euclidean space and by +$\mathbb{R}^{m}_{+}$ the set of all vectors in $\mathbb{R}^{m}$ with nonnegative components. We write $[m]=\{1,\ldots,m\}$. Given a function ${\bm{f}}\colon\mathcal{X}\to\mathbb{R}^{m}$ and a set ${\cal S}\subseteq\mathcal{X}$, we denote by ${\bm{f}}({\cal S})=\{{\bm{f}}(x)\colon x\in{\cal S}\}$ the image of $\cal S$ under $\bm{f}$. Given $x\in\mathcal{X}$ and $r\geq 0$, $B(x,r)=\{y\in\mathcal{X}\colon d(x,y)\leq r\}$ denotes the closed ball centered at $x$ with radius $r$. For a non-empty set ${\cal S}\subseteq\mathbb{R}^{m}$, let $\partial{\cal S}$ denote its boundary. If another non-empty set $\mathcal{S}^{\prime}\subseteq\mathbb{R}^{m}$ is given, then we define the Minkowski sum and difference of $\mathcal{S}$ and $\mathcal{S}^{\prime}$ as $\mathcal{S}+\mathcal{S}^{\prime}=\{\bm{\mu}+\bm{\mu}^{\prime}\colon\bm{\mu}\in\mathcal{S},\bm{\mu}^{\prime}\in\mathcal{S}^{\prime}\}$, $\mathcal{S}-\mathcal{S}^{\prime}=\{\bm{\mu}-\bm{\mu}^{\prime}\colon\bm{\mu}\in\mathcal{S},\bm{\mu}^{\prime}\in\mathcal{S}^{\prime}\}$, respectively. For a vector $\bm{\mu}^{\prime\prime}\in\mathbb{R}^{m}$, we define $\bm{\mu}^{\prime\prime}+\mathcal{S}=\{\bm{\mu}^{\prime\prime}\}+\mathcal{S}$.
@@ 3.1 Multi-objective optimization
Note that the front associated with an $\bm{\epsilon}$-accurate Pareto set is a subset of the $\bm{\epsilon}$-Pareto front. As mentioned in Zuluaga et al. (2016), an $\bm{\epsilon}$-accurate Pareto set is a natural substitute of the Pareto set since any $\bm{\epsilon}$-accurate Pareto design is guaranteed to be no worse than $2\bm{\epsilon}$ of any Pareto optimal design.
3.2 Structure and dimensionality of the design space
In order to learn, we need to make use of some concepts that facilitate understanding the structure of the space and ‘navigating’ it. To that end, we will assume that the design space is well-behaved. This will allow us to partition the design space $\mathcal{X}$ along a tree structure, each level $h$ of which is associated with a partition of $\mathcal{X}$ into $N^{h}$ equal-sized regions $X_{h,i}$, centered at a node $x_{h,i}$, for all $0\leq i\leq N^{h}$, where $N\in\mathbb{N}$. The formal definition is as follows.
Definition 6.
  1. 1.
  2. 2.
  3. 3.
    We assume that the cells have geometrically decaying radii, i.e., there exist $10<\rho<$ 1 and $0<v_{2}\leq 1\leq v_{1}$ such that we have $B(x_{h,i},\,v_{2}\rho^{h})\subseteq X_{h,i}\subseteq B(x_{h,i},\,v_{1}\rho^{h})$ for every $h\geq 0$. Note that we have $2v_{2}\rho^{h}\leq\textup{diam}(X_{h,i})\leq 2v_{1}\rho^{h}$, where $\textup{diam}(X_{h,i})=\textup{sup}_{x,y\in X_{h,i}}d(x,y)$.
[…]
Definition 7.
@@ 4 Adaptive $\epsilon$-PAL algorithm
In this section, we introduce our algorithm, Adaptive $\epsilon$-Pareto Active Learning (PAL), which builds on the $\epsilon$-PAL algorithm of Zuluaga et al. (2016) and utilizes adaptive discretization techniques to enable an efficient navigation of the continuous search space. The algorithm is largely technical, thus we divide the description of each of its phases into more comprehensible subsections.
@@ 4.1 Modeling
Iteration over individual designs may not be feasible when the cardinality of the design space is very large. Thus, we consider partitioning the space into regions of similar designs, i.e., two designs in $\mathcal{X}$ which are at a close distance have similar outcomes in each objective. This is a natural property of GP-sampled functions as discussed in Remark 1. Since our metric space is well-behaved (see Definition 6), there exist an $N\in\mathbb{N}$ and a sequence $(\mathcal{X}_{h})_{h\geq 0}$ of subsets of $\mathcal{X}$ such that for each $h\geq 0$, the set $\mathcal{X}_{h}$ contains $N^{h}$ nodes denoted by $x_{h,1},\ldots,x_{h,N^{h}}$. For each $i\in[N^{h}]$, the associated cell of node $x_{h,i}$ is denoted by $X_{h,i}$.
@@
At round $t\in\mathbb{N}$, the algorithm considers each node $x_{h,i}\in\mathcal{A}_{t}$. Let $j\in[m]$. We define the lower index of $x_{h,i}$ in the $j$th objective as $L^{j}_{t}(x_{h,i})=\underline{B\mkern-4.0mu}\mkern 4.0mu^{j}_{t}(x_{h,i})-V_{h}$, where $\underline{B\mkern-4.0mu}\mkern 4.0mu^{j}_{t}(x_{h,i})$ is a high probability lower bound on the $j$th objective value at $x_{h,i}$ and is defined as
@@
Here, $\beta_{\tau}$ is a parameter to be defined later and $V_{h}$ is a high probability upper bound on the maximum variation of the objective $j$ inside region $X_{h,i}$, which will also be defined later. Similarly, we define the upper index of $x_{h,i}$ in the $j$th objective as $U^{j}_{t}(x_{h,i})=\bar{B}^{j}_{t}(x_{h,i})+V_{h}$, where $\bar{B}^{j}_{t}(x_{h,i})$ is a high probability upper bound on the $j$th objective value at $x_{h,i}$ and is defined as
@@
assuming that $\bm{R}_{t-1}(x_{h,i})$ is well-defined at round $t-1$ (the case $t\geq 2$) or using the convention that $\bm{R}_{0}(x_{0,1})=\mathbb{R}^{m}$ since $\mathcal{A}_{1}=\{x_{0,1}\}$ (the case $t=1$). The well-definedness assumption will be verified in the refining/evaluating phase below.
@@ 4.2 Discarding phase
If a node $x\in\mathcal{A}_{t}$ is $\bm{\epsilon}$-dominated by any other node in $\mathcal{A}_{t}$ under uncertainty, then the algorithm is confident enough to discard it. To check this, the algorithm compares $x$ with all of the pessimistic available points as introduced next.
@@
(Pessimistic Pareto set) Let $t\geq 1$ and let $D\subseteq\mathcal{A}_{t}$ be a set of nodes. We define $p_{pess,t}(D)$, called the pessimistic Pareto set of $D$ at round $t$, as the set of all nodes $x\in D$ for which there is no other node $y\in D\setminus\{x\}$ such that $\textup{min}(\bm{R}_{t}(x))\preceq\textup{min}(\bm{R}_{t}(y))$. We call a design in $p_{pess,t}(D)$ a pessimistic Pareto design of $D$ at round $t$.
@@ 4.3 $\bm{\epsilon}$-Covering phase
For each $x_{h,i}\in\mathcal{S}_{t}$, the algorithm checks if $X_{h,i}$ belongs to an $\mathcal{O}_{\epsilon}$ with high probability in view of Definition 11. In this case, $x_{h,i}$ is removed from the set $\mathcal{S}_{t}$ of undecided nodes and is moved to the set $\mathcal{P}_{t}$ of decided nodes; otherwise, no change is made. The nodes in $\mathcal{P}_{t}$ are never removed from this set; hence, they will be returned by the algorithm as part of the set $\hat{\mathcal{P}}$ at termination. In the appendix, we show that the union $\hat{P}=\bigcup_{x_{h,i}\in\hat{\mathcal{P}}}X_{h,i}$ of the cells is an $\bm{\epsilon}$-accurate Pareto cover, according to Definition 5, with high probability.
Note that while the sets $\mathcal{S}_{t},\mathcal{P}_{t}$ can be modified during this phase, the set $\mathcal{W}_{t}$ does not change.
The modeling, discarding, and $\bm{\epsilon}$-covering phases of the algorithm are illustrated in Figure 1.

### v4

@@ Abstract
Work done while authors were at Bilkent University.
@@ 1 Introduction
We prove information-type and metric dimension-type upper bounds on the sample complexity of Adaptive $\bm{\epsilon}$-PAL (Theorem 1). Our information-type bound yields a sample complexity upper bound of $\tilde{O}(g(\epsilon))$ where $\epsilon=\min_{j}\epsilon^{j}$, $g(\epsilon)=\min\{\tau\geq 1:\sqrt{\gamma_{\tau}/\tau}<\epsilon\}$, and $\gamma_{\tau}$ is the maximum information gain after $\tau$ evaluations. To the best of our knowledge, this is the first information type bound for dependent objectives in the context of PAL. In addition, our metric dimension-type bound yields a sample complexity of $\tilde{O}(\epsilon^{-(\frac{\bar{D}}{\alpha}+2)})$ for all $\bar{D}>D_{1}$, where $D_{1}$ represents the metric dimension of $({\cal X},d)$ and $\alpha\in(0,1]$ represents the Hölder exponent of the metric induced by the GP on $\cal X$}. As far as we know, this is the first metric dimension-type bound in the context of PAL. Our bounds complement each other: as we show in the appendix, neither of them dominates the other for all possible GP kernels. Specifically, we provide an example (Proposition 1) under which the information-type bound can be very loose compared to the metric dimension-type bound. Besides theory, we also show via simulations on multi-objective functions that Adaptive $\bm{\epsilon}$-PAL significantly improves over its competitors in terms of accuracy. Furthermore, we point out the key challenges that necessitate novel algorithmic design and techniques as follows:
  • First, due to the nature of the problem, we need to properly define novel confidence hyper-rectangle objects which are designed to capture uncertainty, not only over individual designs, but over whole regions. This requires integrating information and metric-type components into our hyper-rectangle definitions.
  • Second, the refining, evaluating and discarding steps of our procedure are substantially different from their analogues in $\bm{\epsilon}$-PAL (Zuluaga et al., 2016). Again, this is due to the large space of designs to consider. Here, we need to be confident about regions of space, rather than individual points, and choose whether to discard them or maintain them.
  • Third, in addition to the Pareto front identification problem, inherent in $\bm{\epsilon}$-PAL, our formulation also introduces the problem of efficiently expanding the search tree in promising directions (note that the tree is not uniformly explored). Thus, we need additional analysis which involve information-theoretic and geometric components integrated into a threshold condition.
  • Finally, all of the above necessitate new additional results which involve algebraic and analytic arguments to carry through the analysis. Specifically, results such as Lemma 8 (which requires a careful step-by-step analysis applicable only to our setting), Lemma 10, Lemma 12 and Lemma 16, are all results which require non-trivial and novel reasoning steps necessary for our setting.
Organization. We provide a detailed comparison with related work in Section 2. We explain the properties of the function to be optimized and the structure of the design space in Section 3. This is followed by the description of Adaptive $\bm{\epsilon}$-PAL in Section 4. We give sample complexity bounds for Adaptive $\bm{\epsilon}$-PAL in Section 5, discuss the main aspects of computational complexity analysis in Section 6. We devote Section 7 to the computational experiments, followed by conclusions in Section 8. We present the proof of the main theorem separately in Section B, the appendix. At the end of the paper, we include a table for the frequently used notation.
@@ 2.1 Multi-objective optimization
Auer et al. (2016) consider a finite set of designs and formulate the identification of the Pareto front as a pure exploration multi-armed bandit (MAB) problem in the fixed confidence setting. Under the assumption that the centered outcomes are $1$-subGaussian and independent, they provide gap-dependent bounds on its sample complexity, which, in the single objective case, yield the well-known gap-dependent sample-complexity bounds for pure exploration in the MAB setting (Mannor & Tsitsiklis, 2004). Other works on Multi-objective MAB include (Xu & Klabjan, 2023; Busa-Fekete et al., 2017; Öner et al., 2018). As opposed to their frequentist approach, our approach is Bayesian, which imposes a Gaussian process prior to the latent function of interest.
@@
Ours Compact GP sample Yes Inf. & dim.-type Used
@@ 3 Background and formulation
Throughout the paper, let us fix a positive integer $m\geq 2$ and a compact metric space $(\mathcal{X},d)$. We denote by $\mathbb{R}^{m}$ the $m$-dimensional Euclidean space and by $\mathbb{R}^{m}_{+}$ the set of all vectors in $\mathbb{R}^{m}$ with nonnegative components. We write $m]=\{1,\ldots,m\}$. Given a function ${\bm{f}}\colon\mathcal{X}\to\mathbb{R}^{m}$ and a set ${\cal S}\subseteq\mathcal{X}$, we denote by ${\bm{f}}({\cal S})=\{{\bm{f}}(x)\colon x\in{\cal S}\}$ the image of $\cal S$ under $\bm{f}$. Given $x\in\mathcal{X}$ and $r\geq 0$, $B(x,r)=\{y\in\mathcal{X}\colon d(x,y)\leq r\}$ denotes the closed ball centered at $x$ with radius $r$. For a non-empty set ${\cal S}\subseteq\mathbb{R}^{m}$, let $\partial{\cal S}$ denote its boundary. If another non-empty set $\mathcal{S}^{\prime}\subseteq\mathbb{R}^{m}$ is given, then we define the Minkowski sum and difference of $\mathcal{S}$ and $\mathcal{S}^{\prime}$ as $\mathcal{S}+\mathcal{S}^{\prime}=\{\bm{\mu}+\bm{\mu}^{\prime}\colon\bm{\mu}\in\mathcal{S},\bm{\mu}^{\prime}\in\mathcal{S}^{\prime}\}$, $\mathcal{S}-\mathcal{S}^{\prime}=\{\bm{\mu}-\bm{\mu}^{\prime}\colon\bm{\mu}\in\mathcal{S},\bm{\mu}^{\prime}\in\mathcal{S}^{\prime}\}$, respectively. For a vector $\bm{\mu}^{\prime\prime}\in\mathbb{R}^{m}$, we define $\bm{\mu}^{\prime\prime}+\mathcal{S}=\{\bm{\mu}^{\prime\prime}\}+\mathcal{S}$.
@@ 3.1 Multi-objective optimization
Note that the front associated with an $\bm{\epsilon}$-accurate Pareto set is a subset of the $\bm{\epsilon}$-Pareto front. As mentioned in (Zuluaga et al., 2016), an $\bm{\epsilon}$-accurate Pareto set is a natural substitute of the Pareto set since any $\bm{\epsilon}$-accurate Pareto design is guaranteed to be no worse than $2\bm{\epsilon}$ of any Pareto optimal design.
@@ 4 Adaptive $\epsilon$-PAL algorithm
In this section, we introduce our algorithm, Adaptive $\bm{\epsilon}$-Pareto Active Learning (PAL), which builds on the $\bm{\epsilon}$-PAL algorithm of Zuluaga et al. (2016) and utilizes adaptive discretization techniques to enable an efficient navigation of the continuous search space. The algorithm is largely technical, thus we divide the description of each of its phases into more comprehensible subsections.
Intuition
On a high level, our algorithm’s rationale can be explained as follows. First, Adaptive $\bm{\epsilon}$-PAL maintains an adaptively changing resolution over the design space, structured along so-called nodes of a tree. That is, starting from the center of the space (the center node), we only ‘zoom in’ (thus expanding the tree) on points that are of potential interest. In every iteration of the algorithm, we are given the current set of active nodes in our tree-based partition of the space. These nodes serve as proxies for regions of points in the design space which they inhabit. Now, the goal of the algorithm is to return a set of nodes (and, as a consequence, their associated regions) that form an approximate Pareto front with high probability. In order to do that, the algorithm maintains a confidence region for every active node. Applying worst-case arguments over these confidence regions, the algorithm decides which points to discard in every iteration, and which points to maintain as potentially optimal (in the Pareto sense). Next, we move points about which we are confident enough to a predicted Pareto set. Basically, we keep shrinking the set of points about which we are undecided, and we keep growing the set of points which we believe are approximately optimal. Finally, if enough information is obtained on the most uncertain active node, then we decide to expand it into children nodes, since there is enough reason to believe that more relevant information will be obtained from those new nodes.
@@ 4.1 Modeling
Iteration over individual designs may not be feasible when the cardinality of the design space is very large. Thus, we consider partitioning the space into regions of similar designs, i.e., two designs in $\mathcal{X}$ which are at a close distance have similar outcomes in each objective. This is a natural property of GP-sampled functions as discussed in Remark 1.
Since iteration over individual designs is not possible in large spaces, we will instead iterate over ‘regions’ of interest. Here, we focus on a compact subset $\mathcal{X}$ of the Euclidean space $\mathbb{R}^{m^{\prime}}$ for some $m^{\prime}\in\mathbb{N}$. However, we do this purely for simplicity of presentation. In Appendix, we show that our analysis holds when $\mathcal{X}$ is any general ‘well-behaved’ metric space. [^2] See Section A for detailed definitions. For such a metric space, one can easily partition the design space along a tree structure, each level $h$ of which is associated with a partition of $\mathcal{X}$ into $N^{h}$ equal-sized regions $X_{h,i}$, centered at a node $x_{h,i}$, for all $0\leq i\leq N^{h}$, where $N\in\mathbb{N}$.
@@
At round $t\in\mathbb{N}$, the algorithm considers each active node $x_{h,i}\in\mathcal{A}_{t}$. Let $j\in[m]$. We define the lower index of $x_{h,i}$ in the $j$th objective as $L^{j}_{t}(x_{h,i})=\underline{B\mkern-4.0mu}\mkern 4.0mu^{j}_{t}(x_{h,i})-V_{h}$, where $\underline{B\mkern-4.0mu}\mkern 4.0mu^{j}_{t}(x_{h,i})$ is a high probability lower bound on the $j$th objective value at $x_{h,i}$ and is defined as
@@
Here, $\beta_{\tau}\in O(\log(\tau^{2}/\delta))$ and $V_{h}\in\tilde{O}(\rho^{\alpha h})$ is a high probability upper bound on the maximum variation of the objective $j$ inside region $X_{h,i}$. [^3] See Theorem 1 for exact definitions. Similarly, we define the upper index of $x_{h,i}$ in the $j$th objective as $U^{j}_{t}(x_{h,i})=\bar{B}^{j}_{t}(x_{h,i})+V_{h}$, where $\bar{B}^{j}_{t}(x_{h,i})$ is a high probability upper bound on the $j$th objective value at $x_{h,i}$ and is defined as
@@
assuming that $\bm{R}_{t-1}(x_{h,i})$ is well-defined at round t1t-1 (the case $t\geq 2$) or using the convention that $\bm{R}_{0}(x_{0,1})=\mathbb{R}^{m}$ since $\mathcal{A}_{1}=\{x_{0,1}\}$ (the case $t=1$). The well-definedness assumption will be verified in the refining/evaluating phase below.
@@ 4.2 Discarding phase
If a node $x\in\mathcal{A}_{t}$ is $\bm{\epsilon}$-dominated by any other node in $\mathcal{A}_{t}$ under uncertainty, then the algorithm is confident enough to discard it. Basically, the condition of Definition 7 implies that, if the best possible value that $x$ can have is still approximately dominated by the worst possible value that $y$ can have, then we can conclude that $y$ dominates $x$ with overwhelming probability. To check this, the algorithm compares $x$ with all of the pessimistic available points as introduced next.
@@
(Pessimistic Pareto set) Let $t\geq 1$ and let $D\subseteq\mathcal{A}_{t}$ be a set of nodes. We define $p_{pess,t}(D)$, called the pessimistic Pareto set of $D$ at round $t$, as the set of all nodes $x\in D$ for which there is no other node $y\in D\setminus\{x\}$ such that $\textup{min}(\bm{R}_{t}(x))\prec\textup{min}(\bm{R}_{t}(y))$. We call a design in $p_{pess,t}(D)$ a pessimistic Pareto design of $D$ at round $t$.
@@ 4.3 $\bm{\epsilon}$-Covering phase
For each $x_{h,i}\in\mathcal{S}_{t}$, the algorithm checks if $X_{h,i}$ belongs to an $\mathcal{O}_{\epsilon}$ with high probability in view of Definition 9. In this case, $x_{h,i}$ is removed from the set $\mathcal{S}_{t}$ of undecided nodes and is moved to the set $\mathcal{P}_{t}$ of decided nodes; otherwise, no change is made. The nodes in $\mathcal{P}_{t}$ are never removed from this set; hence, they will be returned by the algorithm as part of the set $\hat{\mathcal{P}}$ at termination. Intuitively, this step determines which points can be safely predicted to be in the approximately accurate Pareto set if there is no other active node which approximately dominates it in the worst case possible. In the appendix, we show that the union $\hat{P}=\bigcup_{x_{h,i}\in\hat{\mathcal{P}}}X_{h,i}$ of the cells is an $\bm{\epsilon}$-accurate Pareto cover, according to Definition 5, with high probability. Note that while the sets $\mathcal{S}_{t},\mathcal{P}_{t}$ can be modified during this phase, the set $\mathcal{W}_{t}$ does not change. The modeling, discarding, and $\bm{\epsilon}$-covering phases of the algorithm are illustrated in Figure 1.
@@ 4.4 Refining/evaluating phase
Figure 4: A depiction of three refining iterations of our algorithm together with the associated discarding and $\bm{\epsilon}$-covering phases. The first row reflects the structural changes in the design space. The second row reflects the changes in the objective space. The green color in the region bars denotes a predicted Pareto region, the orange color denotes an undecided region, while the gray color a discarded region. At time $t$, we have two nodes $x_{1,1}$ and $x_{1,2}$ that partition $\mathcal{X}$. The region $X_{1,2}$ is already a predicted Pareto set, while $X_{1,1}$ is still undecided. This is also reflected in the first figure (from the left) on the second row. Here, we can visually see that $x_{1,1}$ is $\bm{\epsilon}$-dominated by $x_{1,2}$ under uncertainty. Since there are no more nodes, we proceed to assign $X_{1,2}$ to a predicted Pareto set. In the next iteration, we first refine $x_{1,1}$ further because the uncertainty level is already below the threshold. Then, upon examination, we discard region $X_{2,1}$. Again, the corresponding changes in the objective space are reflected in the second row. In the next iteration, we further refine $x_{2,2}$ and include one of its children regions to the predicted Pareto set. Note that we have not included evaluation steps for simplicity of presentation.